# Embedding-based Approaches to Hyperpartisan News Detection

### v1

@@ 1.2 Related Work
Last year, the organizers of SemEval2019 (Kiesel et al., 2019) released a large-scale dataset with over 1M articles in Task 4, Hyper-partisan News Detection, where data are labeled with the tagset hyperpartisan, not-hyperpartisan. Considering that many previous works have not utilized deep learning methods for hyperpartisan detection, one of our goals is to narrow down this gap by exploring how well deep learning can handle this task. We employed feed-forward neural networks and Convolutional Neural Networks (CNN). We also employed various types of features including sentence embeddings, n-grams, and sentiment and emotion features.
The rest of the report is organized as follows: data and task definition are described in Section 2, Section 3 describes our methods, followed by experiments and results in Section 4. Finally, we discuss future works and conclude this report in Section 5.
@@ 2 Data
In this project, we use the data provided by SemEval2019 task 4 (Kiesel et al., 2019). The task is set up as a binary classification problem where news articles are labeled with the target {hyperpartisan, not-hyperpartisan}.
@@
The second, larger dataset (by-publisher corpus) contains 754,000 articles which were automatically labeled based on a categorization of the political bias of news publishers. This dataset was split into a training set of 600,000 articles and a validation set of 150,000 articles. At first, we expected the use of by-publisher training data can help classify the by-article set, but it turned out that though much larger than the by-article set, its labels contain significant amount of noisy introduced by the labeling program (Kiesel et al., 2019). The utilization of the by-publisher set seems more difficult than expected and therefore we decided not to use it.
@@ 3.1.1 N-Grams
The first obvious approach for the task was employing classic ML classifiers using the simple n-gram features. We built a baseline system using unigram and bigram features over a range of term frequency cutoff. The performance peaked when the cutoff is near 10 - 15. We tested a variety of classifiers including Logistic Regression, SVM, Random Forest, and Gradient Boosting Tree.
@@ 3.1.2 Emotionality (Sentiment Analysis)
Valence (sentiment) and polarity analysis is another approach that seems suitable for the task assuming that hyperpartisan news typically involves a high intensity of valence, subjectivity, and polarity because of its biased nature. We used the following features: polarity, subjectivity; and positive, negative and neutral scores, which are obtained from textblob and NLTK text analysis.
@@ 3.1.3 Word Embeddings
Word embeddings are commonly used in many NLP tasks in recent years because they are found to be useful representations of words and often lead to better performance. The input to neural networks is a set of pretrained text representations such as Word2Vec (Mikolov et al., 2013), Glove (Pennington et al., 2014), ELMo (Peters et al., 2018), or BERT (Devlin et al., 2018). In our system we use ELMo embeddings which have the advantage of modeling polysemy and morphological features that word-level embeddings could omit. We represent the article as a sequence of sentence embedding by averaging the word embeddings of that sentence. Considering the range of article size, only the first 250 sentences of an article and the first 250 tokens of a sentence was used for each article representation. We also tried the pretrained BERT for word embedding. The drawback of using BERT is that we were limited to a sequence length of 512, hence we had to restrict the article sequence length to 512 and continue the training process. The Accuracy is 0.81. However this may not be a good embedding model since we are abruptly restricting the article to 512 word embedding.
@@ 3.2 Classifiers and Model Architecture
3.2.1 Feedforward Neural Networks
We first built a system using only fully connected layer with a sequence of sentence embeddings as input as described above. Batch Normalization (Ioffe and Szegedy, 2015) and dropout (Srivastava et al., 2014) was used for better generalization. The network has two hidden fully connected layers with ReLU activation and an output layer with sigmoid activation.
@@ 3.2.3 Long Short Term Memory Networks (LSTMs)
Long Short Term Memory Networks (Hochreiter and Schmidhuber, 1997) have been successfully applied to many text classification problems. RNNs are good summarizer of sequantial information such as language yet suffer from gradient issues. LSTMs solve this issue to some extent. A variation of LSTM is Bidirectional LSTMs (Bi-LSTMs), which summarizes information from both left to right and vice versa. We investigate LSTMs and Bi-LSTMs on this task.
@@ 3.2.4 Other Variety of Classifiers
We also used a variety of traditional machine learning classifiers using n-grams or emotionality scores as input, including Logistic Regression, SVM, Random Forest, Gradient Boosting Tree, and their ensembles.
@@ 4.1 N-gram Models
Table 1 shows the results of our system using unigrams and bigrams with a frequency cutoff 12. Numbers appear in articles are all replaced by a token <num>. Gradient Boosting Trees turns out to be the winner, achieving an 0.775 average accuracy over 10-fold cross validation. In the second place is the ensemble of the four classifiers, indicating that the prediction of all the four models are almost identical so we can’t benefit from ensembling.
@@ 4.2 Sentiment Analysis
Table 2 shows the results of our system using polarity, subjectivity, and valence scores. The results are much worse than n-gram systems. There could be three reasons for the poor performance. First, the methods used to generate the scores were not very optimal. Second, to fully leverage this methods, we should use more hand-crafted lexicon-based features. Finally, we don’t have enough training data to build our own sentiment analysis model.
@@ 4.3 Neural Networks
Table 3 illustrates the average accuracy over 10-fold cross validation obtained on the systems of feedforward neural networks, CNN, and LSTM. Dropout is used for feedforward networks and LSTM and batch normalization is used for feedforward networks and CNN. We used the default Adam as the optimizer and binary cross-entropy as the loss function. We can see a huge improvement by leveraging the complexity of neural networks. The four systems are comparable but Bi-LSTM slightly outperforms the others.
@@ 4.4 Overfitting Investigation
Overfitting has been a problem throughout the entire experiments. This is reasonable considering the size of training size is only 645. Thus, we conducted another experiment on how the number of training examples will influence performance. We randomly chose 80 examples as test set and trained feedforward networks using 50, 100, 200, 350, and 565 examples from the rest of the dataset. Figure 1 illustrates the result, suggesting that our system could perform even better if given a larger training set.
@@ 5 Discussion and Conclusion
In this report, we present our systems on hyperpartisan news detection. We trained a set of models using a variety of features including n-grams, sentiment scores, and word embeddings. The Bi-LSTM system using ELMo embedding acquired a 83.72% accuracy through 10-fold cross validation without much hyperparameter tuning. We also investigated into the overfitting problem and concluded that the potential of deep learning methods has not been fully exploited due to the small training set.
A few things could be done in future work. First, it would be helpful to have a larger sized dataset to handle overfitting issues. Second, we haven’t tuned hyperparameters for each network. Finally, we should try other state-of-the-art models and more expressive network architectures.

### v3

@@ 1.2 Related Work
Last year, the organizers of SemEval2019 (Kiesel et al., 2019) released a large-scale dataset with over 1M articles in Task 4, Hyper-partisan News Detection, where data are labeled with the tagset hyperpartisan, not-hyperpartisan. Considering that many previous works have not utilized deep learning methods for hyperpartisan detection, one of our goals is to narrow down this gap by exploring how well deep learning can handle this task. Feed-forward neural networks and Convolutional Neural Networks (CNN) were using for various types of features including sentence embeddings, n-grams, and sentiment and emotion features.
The rest of the report is organized as follows: data and task definition are described in Section 2, Section 3 describes our methods, followed by experiments and results in Section 4. Finally, I discuss future work and conclude this report in Section 5.
@@ 2 Data
In this report, the data used is provided by SemEval2019 task 4 (Kiesel et al., 2019). The task is set up as a binary classification problem where news articles are labeled with the target {hyperpartisan, not-hyperpartisan}.
@@
The second, larger dataset (by-publisher corpus) contains 754,000 articles which were automatically labeled based on a categorization of the political bias of news publishers. This dataset was split into a training set of 600,000 articles and a validation set of 150,000 articles. At first, I expected the use of by-publisher training data can help classify the by-article set, but it turned out that though much larger than the by-article set, its labels contain significant amount of noisy introduced by the labeling program (Kiesel et al., 2019). The utilization of the by-publisher set seems more difficult than expected and therefore I decided not to use it.
@@ 3.1.1 N-Grams
The first obvious approach for the task was employing classic ML classifiers using the simple n-gram features. A baseline system is built using unigram and bigram features over a range of term frequency cutoff. The performance peaked when the cutoff is near 10 - 15. A variety of classifiers were tested including Logistic Regression, SVM, Random Forest, and Gradient Boosting Tree.
@@ 3.1.2 Emotionality (Sentiment Analysis)
Valence (sentiment) and polarity analysis is another approach that seems suitable for the task assuming that hyperpartisan news typically involves a high intensity of valence, subjectivity, and polarity because of its biased nature. The following features were used: polarity, subjectivity; and positive, negative and neutral scores, which are obtained from textblob and NLTK text analysis.
@@ 3.1.3 Word Embeddings
Word embeddings are commonly used in many NLP tasks in recent years because they are found to be useful representations of words and often lead to better performance. The input to neural networks is a set of pretrained text representations such as Word2Vec (Mikolov et al., 2013), Glove (Pennington et al., 2014), ELMo (Peters et al., 2018), or BERT (Devlin et al., 2018). In our system ELMo embeddings were used which have the advantage of modeling polysemy and morphological features that word-level embeddings could omit. Further, articles are represented as a sequence of sentence embeddings by averaging the word embeddings of those sentences. Considering the range of article size, only the first 250 sentences of an article and the first 250 tokens of a sentence was used for each article representation to account for the 512 sequence limit with pretrained BERT models. The Accuracy is 0.81. However, this may not be a good embedding model due to restriction of the sequence length.
@@ 3.2.2 Long Short Term Memory Networks (LSTMs)
Long Short Term Memory Networks (Hochreiter and Schmidhuber, 1997) have been successfully applied to many text classification problems. RNNs are good summarizer of sequantial information such as language yet suffer from gradient issues. LSTMs solve this issue to some extent. A variation of LSTM is Bidirectional LSTMs (Bi-LSTMs), which summarizes information from both left to right and vice versa. The accuracy scores for LSTMs and Bi-LSTMs are available in table 3.
@@ 3.2.3 Other Variety of Classifiers
A variety of traditional machine learning classifiers using n-grams or emotionality scores as input, including Logistic Regression, SVM, Random Forest, Gradient Boosting Tree, and their ensembles were used.
@@ 4.1 N-gram Models
Table 1 shows the results of our system using unigrams and bigrams with a frequency cutoff 12. Numbers appear in articles are all replaced by a token <num>. Gradient Boosting Trees turns out to be the winner, achieving an 0.775 average accuracy over 10-fold cross validation. In the second place is the ensemble of the four classifiers, indicating that the prediction of all the four models are almost identical and there is no benefit from ensembling.
@@ 4.2 Sentiment Analysis
Table 2 shows the results of our system using polarity, subjectivity, and valence scores. The results are much worse than n-gram systems. There could be three reasons for the poor performance. First, the methods used to generate the scores were not very optimal. Second, to fully leverage this methods, more hand-crafted lexicon-based features were used.
@@ 4.3 Neural Networks
Table 3 illustrates the average accuracy over 10-fold cross validation obtained on the systems of CNN and LSTMs. Dropout is used for feedforward networks and LSTM and batch normalization is used for feedforward networks and CNN. The four systems are comparable but LLM significantly outperforms the others.
@@ 4.4 Overfitting Investigation
Overfitting has been a problem throughout the experiments. This is reasonable considering that the size of training size is only 645. Thus, experiments were conducted on how the number of training examples will influence performance. Figure 1 illustrates the result, suggesting that the system could perform even better if given a larger training set.
@@
4.5 LLMs
Recently, LLMs are being used to generate embeddings. These embeddings capture more nuanced representation in identifying hyperpartisan news. Unlike traditional word embeddings, such as Word2Vec or GloVe, which rely primarily on local context and static representations, embeddings generated by large language models incorporate dynamic contextual information. This allows for better detection of subtle linguistic cues, ideological framing, and rhetorical strategies commonly present in hyperpartisan content. As a result, these embeddings improve the performance of downstream classification models, enhancing their ability to distinguish between balanced reporting and biased narratives.
Additionally, leveraging embeddings from LLMs enables transfer learning across datasets, making models more robust in real-world applications where labeled partisan data may be scarce.
The embeddings generated using Llama-4-Scout-17B-16E-Instruct improved the classifier accuracy to 92.62% while the embeddings generated using gemma3 models improved the classifier accuracy to 91.44%
@@ 5 Discussion and Conclusion
In this report, I present the systems on hyperpartisan news detection. I trained a set of models using a variety of features including n-grams, sentiment scores, and word embeddings. The Bi-LSTM system using ELMo embedding acquired a 83.72% accuracy through 10-fold cross validation without much hyperparameter tuning. The overfitting problem investigation concludes that potential of deep learning methods has not been fully exploited due to the small training set. Lastly, the embeddings generated using LLMs achieved an accuracy of 92.62%. Embeddings generated using LLMs improved the classifier accuracy over the previously built system using ELMo models.